OpenStudy (anonymous):

Write the equation of the line that passes through the points (1, 7) and (5, 15) using function notation. y = 2x + 5 y = 4x + 8 f(x) = 2x + 5 f(x) = 4x + 8

Mehek (mehek14):

function notion = f(x) so not A or B

Mehek (mehek14):

now find the slope \(\dfrac{15-7}{5-1}=\dfrac{8}{4}=2\)

Mehek (mehek14):

so C is your answer

OpenStudy (anonymous):

Thanks so much, you got yourself a medal :)

Mehek (mehek14):

your welcome and thank you very much \(\ddot\smile\)